OpenStudy (anonymous):

write and equation for the nth term of each geometric sequence -1,-5,-25..... and -3,-6,-12,....

OpenStudy (anonymous):

1) -125 2) -24

OpenStudy (anonymous):

Now award best response

OpenStudy (anonymous):

lolz tank @gamer456148

OpenStudy (jmark):

The nth term of Geometric sequence is given by Tn = ar^(n-1) for the sequence -1,-5,-25.... a = -1, r = 5 nth term of sequence is Tn = (-1)x5^(n-1) = -5^(n-1) -3,-6,-12,.... a = -3, r = 2 nth term of sequence is Tn = (-3)x2^(n-1) = -6^(n-1) maths question paper for class 7 ICSE - http://icse.edurite.com/icse-question-papers/icse-class-7-maths-question-papers.html